OpenStudy (anonymous):

HELP PLEASE: Show all steps. solve the system by substitution. -x-y-z=-8 -4x-4y+5z=7 2x+2z=4

OpenStudy (campbell_st):

divide the last equation by 2 and you get x + z = 2 looking at the 1st equation you get -(x + z) -y = -8 substituting and you get -(2) -y = -8 solve for y

OpenStudy (anonymous):

Do I have to get Y alone in order to solve for Y?

OpenStudy (campbell_st):

yep... just add 2 to both sides of the equation 1st then divide both sides by -1

OpenStudy (anonymous):

y=4?

OpenStudy (campbell_st):

well I thought y = 6 but the next step is to substitute y into equation 2 -4x -24 + 5z = 7 or -4x + 5z = 31

OpenStudy (anonymous):

Yeah it would be 6 I did my math wrong. i was dividing -8 by 2 instead of adding 2 so it would be 6

OpenStudy (campbell_st):

ok... so that the last equation again make z the subject 2z = 4 - 2x divide by 2 z = 4 - 2x substitute it into the equation -4x + 5(4 - 2x) = 31 and solve for x

OpenStudy (anonymous):

-6x+9=31 subtract 9 from both sides then divide by -6 to get -3.7 i feel like i did this all wrong?

OpenStudy (campbell_st):

ok... well just checking it... it does have factional answers... and y = 6 is correct I just realised a mistake if you substitute you are solving -4x +5(2 - x) = 31 now -4x + 10 - 5x = 31 or -9x + 10 = 31 -9x = 21 x = -7/3